# Quantum Wave Simulation with Sources and Loss Functions

### v1

@@
Quantum Wave Simulation with Sources and Loss Functions
Cyrill Bösch,1,2,† Malte Schade,1,† Giacomo Aloisi,1 Andreas Fichtner1
1 Department of Earth and Planetary Sciences
   ETH Zurich    8092 Zurich    Switzerland.
2 Department of Computer Science
   Princeton University    Princeton    NJ 08540    USA
Email: cb7454@princeton.edu
These authors contributed equally
keywords:
Numerical modeling, Wave propagation, Tomography, Quantum computing, Quantum algorithms, Quantum simulation
{summary}
@@ 2 Department of Computer Science
We present a quantum algorithmic framework for simulating linear, anti-Hermitian (lossless) wave equations in heterogeneous, anisotropic, but time-independent media. This framework encompasses a broad class of wave equations, including the acoustic wave equation, Maxwell’s equations, and the elastic wave equation. Our formulation is compatible with standard numerical discretization schemes and allows for efficient implementation of source terms for diverse time- and space-dependent sources. Furthermore, we demonstrate that subspace energies can be extracted and wave fields compared through an $l_{2}$-loss function, achieving optimal precision scaling with the number of samples taken. Additionally, we introduce techniques for incorporating boundary conditions and linear constraints that preserve the anti-Hermitian nature of the equations. Leveraging the Hamiltonian simulation algorithm, our framework achieves a quartic speed-up over classical solvers in 3D simulations, under conditions of sufficiently global measurements and compactly supported sources and initial conditions. We argue that this quartic speed-up is optimal for time domain solutions, as the Hamiltonian of the discretized wave equations has local couplings. In summary, our framework provides a versatile approach for simulating wave equations on quantum computers, offering substantial speed-ups over state-of-the-art classical methods.
1 Introduction
@@
Wave-based inverse problems are prevalent in disciplines such as seismology [1, 2, 3, 4], medical imaging [5, 6, 7], non-destructive testing [8, 9, 10], and meta-material research [11, 12, 13, 14]. However, the advancement of these fields is fundamentally limited by the current state of computational resources. In particular, the resolution and uncertainty quantification within imaging problems remain heavily restricted by computational capacities [15, 16, 17, 18, 19, 20, 21].
@@
Quantum Computers (QCs) offer promising runtime improvements for numerous computational problems [22, 23, 24, 25]. One of the most promising groups of quantum algorithms is Hamiltonian Simulation (HS) algorithms. HS provides a framework for simulating quantum systems, i.e., the evolution of the Schrödinger equation, with up to exponential runtime advantages over classical solvers [26, 27, 28, 29]. A quantum state represented by $n$ two-state quantum particles or qubits has a state space dimension of $N=2^{n}$. Classical algorithms for simulating its evolution under a quantum Hamiltonian $\mathbf{H}\in\mathbb{C}^{2^{n}\times 2^{n}}$ scale at least linearly with the state size $N$ and, hence, exponentially with the number of qubits $n$. In contrast, under specific conditions utilizing HS algorithms, a Quantum Computer (QC) is capable of performing quantum simulations with logarithmic scaling, thereby achieving an exponential speed-up compared to all classical computers:
Theorem 1 (Optimal sparse Hamiltonian simulation, ad. [27] - Theorem 3).
@@
A $d$-sparse Hamiltonian H on $n$ qubits with matrix elements specified to $m$ bits of precision can be simulated for time-interval $t$, error $\epsilon$, and success probability at least $1-2\epsilon$ with $\mathcal{O}(td||\textbf{H}||_{\text{max}}+\frac{\log(1/\epsilon)}{\log\>\log(1/\epsilon)})$ queries to an oracle $\mathcal{H}$ that provides a description of $\mathbf{H}$ and a factor $\mathcal{O}(n+m\>\text{polylog}(m))$ additional quantum gates.
@@
The apparent question is: Can HS algorithms be used to simulate classical wave equations with the same runtime advantages? A growing body of research has explored the potential of HS for classical wave propagation problems. These studies examine source-free wave equations in homogeneous media [30, 31] or wave equations in media with varying parameters [32, 33, 34]. Each of these studies is focusing on a specific numerical approach, predominantly employing Finite Difference (FD) methods. Additionally, recent research has investigated solutions to mass-spring models [35, 36] that yield exponential runtime advantages in certain settings, with particular types of source terms considered in [37], or wave simulation problems addressed through warped phase transformations [38, 39, 40].
In this work, we demonstrate that for loss-free wave equations, including the elastic, acoustic, and electromagnetic case, a natural quantum encoding can be employed (Section 2), which enables a direct measurement of the wave field energy. Furthermore, by introducing the quantum encoding in the continuous domain, we effectively decouple it from discretization requirements, thus rendering our framework compatible with any numerical method.
We proceed by introducing algorithms to measure $l_{2}$-related norms and subspace energies for extracting information from the simulated wave field (Section 3). Additionally, we propose a method to map sources to initial conditions, allowing efficient initialization on the QC (Section 4). Our overall objective is to establish an end-to-end speedup encompassing source initialization, simulation, and final state measurement. Achieving this requires each component to function cohesively, thereby realizing an advantage over classical computation.
Our primary result demonstrates that for sources and initial conditions with compact support, and when the measured quantities of the final state have sufficiently global characteristics relative to the system size, the proposed algorithmic framework achieves a quartic speedup in the three-dimensional case compared to all classical wave equation solvers. We argue that this runtime-advantage is optimal for time domain solutions of wave equations because of local couplings (Section 5).
@@
2 A Natural Quantum Encoding for Linear, Anisotropic, Heterogeneous but Lossless Wave Equations
We consider wave equations of the form
$\frac{d\hat{\mathbf{w}}(\mathbf{x},t)}{dt}=\hat{\mathbf{C}}(\mathbf{x})\hat{\mathbf{w}}(\mathbf{x},t)+\hat{\mathbf{s}}(\mathbf{x},t),\quad\hat{\mathbf{w}}(\mathbf{x},0)=\hat{\mathbf{w}}_{0}(\mathbf{x}),\quad\hat{\mathbf{P}}_{\partial\Omega}\hat{\mathbf{w}}(\overline{\mathbf{x}},t)=\mathbf{0},\forall\overline{\mathbf{x}}\in\partial\Omega$ (1)
@@ 2 A Natural Quantum Encoding for Linear, Anisotropic, Heterogeneous but Lossless Wave Equations
where the wave field $\hat{\mathbf{w}}$ is an element of a Hilbert space $\mathbb{H}^{L}$ containing at least once-differentiable and square-integrable $L$-dimensional vector fields and $\hat{\mathbf{s}}$ is the source term. For concreteness, we have chosen Dirichlet Boundary Conditions on the domain boundary $\partial\Omega$, where $\hat{\mathbf{P}}_{\partial\Omega}\in\mathbb{R}^{L\times L}$ is a diagonal projector which selects the elements of the wave field on which we impose homogeneous Dirichlet BCs.
The operator $\hat{\mathbf{C}}$ is defined as
@@
where $\hat{\mathbf{B}}:\mathbb{H}^{L}\rightarrow\mathbb{H}^{L}$ is any Hermitian (self-adjoint) and positive-definite operator, and $\hat{\mathbf{A}}:\mathbb{H}^{L}\rightarrow\mathbb{H}^{L}$ is any anti-Hermitian (anti-self-adjoint) operator with respect to the inner product
@@
where $\braket{\hat{\mathbf{w}}}{\hat{\mathbf{w}}^{\prime}}=\sum_{l=1}^{L}\hat{w}_{l}^{*}\hat{w}^{\prime}_{l}$, * denotes complex conjugation, and $\hat{\mathbf{w}}\in\mathbb{H}^{L}$ and $\hat{\mathbf{w}}^{\prime}\in\mathbb{H}^{L}$ are any two wave fields. The operator $\hat{\mathbf{C}}$ is therefore anti-Hermitian with respect to the inner product
@@
$\braket{\hat{\mathbf{w}}(\mathbf{x},t)}{\hat{\mathbf{B}}(\mathbf{x})}{\hat{\mathbf{w}}^{\prime}(\mathbf{x},t)}=\sum_{l=1}^{L}\sum_{m=1}^{L}\hat{w}_{l}^{*}(\mathbf{x},t)\hat{B}_{l,m}(\mathbf{x})\hat{w}^{\prime}_{m}(\mathbf{x},t).$ (5)
As we will discuss shortly, the inner product in (4) gives rise to a notion of energy that can be estimated on a QC. The anti-Hermiticity of $\hat{\mathbf{C}}$ with respect to $(\cdot|\cdot)_{\hat{\mathbf{B}}}$ is formalized by showing that
@@
using integration by parts and invoking the BCs [41, 34]. Quantum states are naturally normalized with respect to a norm induced by the “quantum inner product” depicted in (3). Therefore, mapping (1) to a Schrödinger equation amounts to transforming $\hat{\mathbf{C}}\rightarrow\hat{\mathbf{C}}_{Q}$, such that $\hat{\mathbf{C}}_{Q}$ is anti-Hermitian with respect to $(\cdot|\cdot)_{\mathds{1}}$. Because $\hat{\mathbf{B}}$ is Hermitian and positive-definite, it has a unique square root $\hat{\mathbf{B}}^{1/2}$, which is also Hermitian and positive-definite. Transforming (1) with $\hat{\mathbf{w}}_{Q}:=\hat{\mathbf{B}}^{1/2}\hat{\mathbf{w}}$ and $\hat{\mathbf{C}}_{Q}:=\hat{\mathbf{B}}^{-1/2}\hat{\mathbf{A}}\hat{\mathbf{B}}^{-1/2}$ leads to
$\frac{d\hat{\mathbf{w}}_{Q}(\mathbf{x},t)}{dt}=\hat{\mathbf{B}}^{-1/2}(\mathbf{x})\hat{\mathbf{A}}(\mathbf{x})\hat{\mathbf{B}}^{-1/2}(\mathbf{x})\hat{\mathbf{w}}_{Q}(\mathbf{x},t),\quad\hat{\mathbf{w}}_{Q}(\mathbf{x},0)=\hat{\mathbf{B}}^{1/2}(\mathbf{x})\hat{\mathbf{w}}_{0}(\mathbf{x}),\quad\hat{\mathbf{B}}^{1/2}(\overline{\mathbf{x}})\hat{\mathbf{P}}_{\partial\Omega}\hat{\mathbf{w}}(\overline{\mathbf{x}},t)=\mathbf{0},\forall\overline{\mathbf{x}}\in\partial\Omega.$ (7)
We ignore the source term here, as it can be incorporated into the initial conditions or handled separately, as we will show in Section 4. Clearly, because $\hat{\mathbf{A}}$ is anti-Hermitian with respect to $(\cdot|\cdot)_{\mathds{1}}$, it follows that $\hat{\mathbf{C}}_{Q}=\hat{\mathbf{B}}^{-1/2}\hat{\mathbf{A}}\hat{\mathbf{B}}^{-1/2}$ is also anti-Hermitian with respect to the same “quantum inner product”. Multiplying (7) by the imaginary unit $i$, we obtain a Schrödinger equation in its usual form with the corresponding Hamiltonian
@@
Finally, observe that
@@
2.1 The Acoustic Wave Equation
For $D$-dimensional acoustics, we have that $\hat{\mathbf{w}}_{\text{acoustic}}=[u;\mathbf{v}]\in\mathbb{R}^{D+1}$, where $u\in\mathbb{R}$ is the scalar pressure field and $\mathbf{v}\in\mathbf{R}^{D}$ is the velocity field. The respective operators are given by [42]
$\hat{\mathbf{B}}_{\text{acoustic}}=\begin{bmatrix}\frac{1}{\rho(\mathbf{x})c^{2}(\mathbf{x})}&\mathbf{0}_{1\times D}\\ \mathbf{0}_{D\times 1}&\rho(\mathbf{x})\mathds{1}_{D\times D}\end{bmatrix},$ (10)
and
@@ 2.1 The Acoustic Wave Equation
2.2 Maxwell’s Equations
For Maxwell’s equations in $D$ dimensions, the state vector is defined as $\hat{\mathbf{w}}_{\text{Maxwell}}=[\mathbf{E};\mathbf{H}]\in\mathbb{C}^{2D}$, where $\mathbf{E}$ represents the electric field and $\mathbf{H}$ represents the magnetic field. The operator containing the material properties is given by [41, 43]
$\hat{\mathbf{B}}_{\text{Maxwell}}=\begin{bmatrix}\boldsymbol{\epsilon}(\mathbf{x})&\mathbf{0}_{D\times D}\\ \mathbf{0}_{D\times D}&\boldsymbol{\mu}(\mathbf{x})\end{bmatrix},$ (13)
where $\boldsymbol{\epsilon}\in\mathbb{C}^{D\times D}$ is the dielectric permittivity tensor and $\boldsymbol{\mu}\in\mathbb{C}^{D\times D}$ is the magnetic permeability tensor. These tensors reduce to a single coefficient times the identity for isotropic materials. For the presented framework to be applicable, both must be positive definite and Hermitian, thereby specifying a generally anisotropic but transparent medium without any losses [44].

### v3

# thanks: These two authors contributed equallythanks: These two authors contributed equally

@@
Abstract
@@
We present a quantum algorithmic framework for simulating linear, anti-Hermitian (lossless) wave equations in heterogeneous, anisotropic, and time-independent media. This framework encompasses a broad class of wave equations, including the acoustic wave equation, Maxwell’s equations and the elastic wave equation. Our formulation is compatible with standard numerical discretization schemes and allows for the efficient implementation of multiple practically relevant time- and space-dependent sources. Furthermore, we demonstrate that subspace energies can be extracted and wave fields compared through an $l_{2}$ loss function, achieving optimal precision scaling with the number of samples taken. Additionally, we introduce techniques for incorporating boundary conditions and linear constraints that preserve the anti-Hermitian nature of the equations. Leveraging the Hamiltonian simulation algorithm, our framework achieves a quartic speed-up over classical solvers in 3D simulations, under conditions of sufficiently global measurements and compactly supported sources and initial conditions. This quartic speed-up is optimal for time-domain solutions, as the Hamiltonian of the discretized wave equations has local couplings. In summary, our framework provides a versatile approach for simulating wave equations on quantum computers, offering substantial speed-ups over state-of-the-art classical methods.
preprint: APS/123-QED
I Introduction
@@
Wave-based inverse problems are prevalent in disciplines such as seismology [1, 2, 3, 4], medical imaging [5, 6, 7], non-destructive testing [8, 9, 10] and metamaterial research [11, 12, 13, 14]. However, these fields are fundamentally limited by the current state of computational resources due to the excessive computational cost of the numerical wave simulation. In particular, resolution and uncertainty quantification in imaging problems remain heavily restricted by computational capacities [15, 16, 17, 18, 19, 20, 21].
@@
Quantum computers (QCs) offer promising runtime improvements for numerous computational problems [22, 23, 24, 25]. One of the most promising groups of quantum algorithms is Hamiltonian Simulation (HS). It provides a framework for simulating quantum systems, i.e., the evolution of the Schrödinger equation, with up to exponential runtime advantages over classical solvers [26, 27, 28, 29]. A quantum state represented by $n$ two-state quantum particles or qubits has a state space dimension of $N=2^{n}$. Classical algorithms for simulating the evolution of a quantum state under a quantum Hamiltonian $\mathbf{H}\in\mathbb{C}^{N\times N}$ scale at least linearly with the state size $N$. In contrast, under specific conditions utilizing HS algorithms, a QC is capable of performing quantum simulations with logarithmic scaling, thereby achieving an exponential speed-up compared to all classical computers. This can be formalized as:
Theorem 1 (Optimal sparse Hamiltonian simulation, adapted [27] - Theorem 3).
@@
A $d$-sparse Hamiltonian $\mathbf{H}$ acting on $n$ qubits with matrix elements specified to $m$ bits of precision can be simulated for time-interval $t$, $l_{2}$-error $\epsilon$, and success probability at least $1-2\epsilon$ with $\mathcal{O}(td||\mathbf{H}||_{\text{max}}+\frac{\log(1/\epsilon)}{\log\>\log(1/\epsilon)})$ queries to an oracle $\mathcal{H}$ that provides a description of $\mathbf{H}$ and a factor $\mathcal{O}(n+m\>\text{polylog}(m))$ additional quantum gates.
@@
The apparent question is: Can HS algorithms be used to simulate classical wave equations with the same runtime advantages? A growing body of research answers this question affirmatively. These studies examine source-free wave equations in homogeneous media [30, 31] or wave equations in media with varying parameters [32, 33, 34, 35]. Furthermore, recent research has investigated solutions to mass-spring models [36, 37] that yield exponential runtime advantages in certain settings, with particular types of source terms considered in [38], or wave simulation problems addressed through warped phase transformations [39, 40, 41].
While many existing studies tend to focus on simulation tasks, our research expands the scope to address the complete end-to-end stack by incorporating a wide range of practically relevant source scenarios and loss function measurements. We thoroughly analyze the conditions necessary to preserve quantum speedup throughout the entire algorithmic stack. Our primary result is to demonstrate that for many practically relevant sources that are either compact or rotationally symmetric, and when the measured quantities of the final state have sufficiently global characteristics relative to the system size, the proposed algorithmic framework achieves a quartic speed-up in the three-dimensional case compared to all classical wave equation solvers.
The rest of the paper is organized as follows. In Section II, we demonstrate that for lossless wave equations, including the elastic, acoustic, and electromagnetic case, a natural quantum encoding can be employed, which enables a direct measurement of the wave field energy and $l_{2}$-norms. We then analyze the runtime-advantage for time-domain solutions of wave equations (Section III). In Section IV we introducing algorithms to measure $l_{2}$-related norms and subspace energies for extracting information from the simulated wave field (Section IV). Then we propose a method for the implementation of sources on a QC (Section V). Finally, we analyze the limitations of our framework (Section VI) and draw conclusions (Section VIII).
@@
II A Natural Quantum Encoding for Linear, Anisotropic, Heterogeneous but Lossless Wave Equations
We consider continuous wave equations of the form
$\frac{d\hat{\mathbf{w}}(\mathbf{x},t)}{dt}$ $=\hat{\mathbf{C}}(\mathbf{x})\hat{\mathbf{w}}(\mathbf{x},t)+\hat{\mathbf{s}}(\mathbf{x},t),$ (1)
$\hat{\mathbf{w}}(\mathbf{x},0)$ $=\hat{\mathbf{w}}_{0}(\mathbf{x}),$
$\hat{\mathbf{P}}_{\partial\Omega}\hat{\mathbf{w}}(\mathbf{x},t)$ $=\mathbf{0},\forall\mathbf{x}\in\partial\Omega,$
@@ II A Natural Quantum Encoding for Linear, Anisotropic, Heterogeneous but Lossless Wave Equations
where hats ( $\hat{}$ ) denote continuous fields and operators over $D$-dimensional space, $\mathbf{x}\in\mathbb{R}^{D}$, and time, $t$. The wave field $\hat{\mathbf{w}}$ is an element of a Hilbert space =$\mathcal{H}=\{f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{C}\}$ containing at least once-differentiable (in both space and time) and square-integrable $C$-dimensional vector fields over the domain $\Omega$, and $\hat{\mathbf{s}}$ is the source term. For concreteness and without loss of generality, we have chosen Dirichlet boundary conditions (BC) on the domain boundary $\partial\Omega$, where $\hat{\mathbf{P}}_{\partial\Omega}\in\mathcal{B}(\mathcal{H})$ is a diagonal projector, which selects the elements of the wave field on which we impose homogeneous Dirichlet BC.
The wave operator $\hat{\mathbf{C}}$ considered here is assumed to have the following form
@@
where $\hat{\mathbf{B}}:\mathcal{H}\rightarrow\mathcal{H}$ is any Hermitian (self-adjoint) and positive-definite operator, and $\hat{\mathbf{A}}:\mathcal{H}\rightarrow\mathcal{H}$ is any anti-Hermitian (anti-self-adjoint) operator with respect to the inner product
@@
where $\braket{\hat{\mathbf{w}}}{\hat{\mathbf{w}}^{\prime}}=\sum_{l=1}^{L}\hat{w}_{l}^{*}\hat{w}^{\prime}_{l}$, (*) denotes complex conjugation, and $\hat{\mathbf{w}},\hat{\mathbf{w}}^{\prime}\in\mathcal{H}$ are any two vector fields. Here, $\hat{\mathbf{A}}$ incorporates various types of spatial derivatives such as divergence, gradient or curl operations. The operator $\hat{\mathbf{C}}$ is therefore anti-Hermitian with respect to the inner product
@@
$\braket{\hat{\mathbf{w}}(\mathbf{x},t)}{\hat{\mathbf{B}}(\mathbf{x})}{\hat{\mathbf{w}}^{\prime}(\mathbf{x},t)}$ (5)
$=\sum_{l=1}^{L}\sum_{m=1}^{L}\hat{w}_{l}^{*}(\mathbf{x},t)\hat{B}_{l,m}(\mathbf{x})\hat{w}^{\prime}_{m}(\mathbf{x},t).$
As we will discuss shortly, the inner product in Eq. (4) gives rise to a notion of energy that can be estimated on a QC. The anti-Hermiticity of $\hat{\mathbf{C}}$ with respect to $(\cdot|\cdot)_{\hat{\mathbf{B}}}$ is formalized by showing that
@@
using integration by parts and invoking the BC [42, 35]. Quantum states are naturally normalized with respect to a norm induced by the inner product (3). Therefore, mapping Eq. (1) to a Schrödinger equation amounts to transforming $\hat{\mathbf{C}}\rightarrow\hat{\mathbf{C}}_{Q}$, such that $\hat{\mathbf{C}}_{Q}$ is anti-Hermitian with respect to $(\cdot|\cdot)_{\mathds{1}}$ [43, 32]. Because $\hat{\mathbf{B}}$ is Hermitian and positive-definite, it has a unique square root $\hat{\mathbf{B}}^{1/2}$, which is also Hermitian and positive-definite. Transforming Eq. (1) with $\hat{\mathbf{w}}_{Q}:=\hat{\mathbf{B}}^{1/2}\hat{\mathbf{w}}$ and $\hat{\mathbf{C}}_{Q}:=\hat{\mathbf{B}}^{-1/2}\hat{\mathbf{A}}\hat{\mathbf{B}}^{-1/2}$ leads to
$\frac{d\hat{\mathbf{w}}_{Q}(\mathbf{x},t)}{dt}$ $=\hat{\mathbf{B}}^{-1/2}(\mathbf{x})\hat{\mathbf{A}}(\mathbf{x})\hat{\mathbf{B}}^{-1/2}(\mathbf{x})\hat{\mathbf{w}}_{Q}(\mathbf{x},t),$ (7)
$\hat{\mathbf{w}}_{Q}(\mathbf{x},0)$ $=\hat{\mathbf{B}}^{1/2}(\mathbf{x})\hat{\mathbf{w}}_{0}(\mathbf{x}),$
$\hat{\mathbf{B}}^{1/2}(\mathbf{x})$ $\hat{\mathbf{P}}_{\partial\Omega}\hat{\mathbf{w}}(\mathbf{x},t)=\mathbf{0},\,\forall\mathbf{x}\in\partial\Omega.$
We ignore the source term here, as it can be incorporated into the initial conditions or handled separately, as we will show in Section V. Clearly, because $\hat{\mathbf{A}}$ is anti-Hermitian with respect to $(\cdot|\cdot)_{\mathds{1}}$, it follows that $\hat{\mathbf{C}}_{Q}=\hat{\mathbf{B}}^{-1/2}\hat{\mathbf{A}}\hat{\mathbf{B}}^{-1/2}$ is also anti-Hermitian with respect to the same inner product. Multiplying Eq. (7) by the imaginary unit $i$, we obtain a Schrödinger equation with the corresponding Hamiltonian
@@
Finally, we observe that
@@
II.1 The Acoustic Wave Equation
For acoustics in $D$ space dimensions, we have that $\hat{\mathbf{w}}_{\text{acoustic}}=[\hat{u};\hat{\mathbf{v}}]\in\{f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D+1}\}$, where $\hat{u}$ is the scalar pressure field and $\hat{\mathbf{v}}$ is the particle velocity field. The respective operators are given by [44]
$\hat{\mathbf{B}}_{\text{acoustic}}(\mathbf{x})=\begin{bmatrix}\frac{1}{\rho(\mathbf{x})c^{2}(\mathbf{x})}&\mathbf{0}_{1\times D}\\ \mathbf{0}_{D\times 1}&\rho(\mathbf{x})\mathds{1}_{D\times D}\end{bmatrix},$ (10)
where $c,\rho\in\{f:\mathbb{R}^{D}\rightarrow\mathbb{R}\}$ are the wave speed and density, respectively. The anti-Hermitian operator is given by
@@ II.1 The Acoustic Wave Equation
[…]
$=2E_{\text{potential}}(t)+2E_{\text{kinetic}}(t).$
II.2 Maxwell’s Equations
For Maxwell’s equations in $D$ dimensions, the state vector is defined as $\hat{\mathbf{w}}_{\text{Maxwell}}=[\hat{\mathbf{E}};\hat{\mathbf{H}}]\in\{f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{2D}\}$, where $\hat{\mathbf{E}}$ represents the electric field and $\hat{\mathbf{H}}$ represents the magnetic field. The operator containing the material properties is given by [42, 45]
$\hat{\mathbf{B}}_{\text{Maxwell}}(\mathbf{x})=\begin{bmatrix}\boldsymbol{\epsilon}(\mathbf{x})&\mathbf{0}_{D\times D}\\ \mathbf{0}_{D\times D}&\boldsymbol{\mu}(\mathbf{x})\end{bmatrix},$ (13)
where $\boldsymbol{\epsilon}$ and $\boldsymbol{\mu}$ are the generally complex but Hermitian dielectric permittivity and magnetic permeability tensors, respectively. These tensors reduce to a single coefficient multiplied with the identity for isotropic materials. For the presented framework to be applicable, both must be positive definite and Hermitian, thereby specifying a generally anisotropic but transparent medium without any losses [46].